OpenStudy (anonymous):

Show that the wave fucntion y=e^b(x-vt) is solution to linear wave equation(i.e d2y/dx2=1/v^2 d2y/dt2? where d is partial derivative

OpenStudy (irishboy123):

just plug it in, eg \(y_t = -vb e^{b(x-vt)}\), and then add it all up...

OpenStudy (anonymous):

plz explain with steps sir if possible

OpenStudy (anonymous):

i dont know much about derivatives and integrals

OpenStudy (irishboy123):

lol! seriously, if you don't know that *much about* calculus, messing about with the wave equation is a tough one for anyone :p you have \(y_t\) which to be clear is a simpler way of writing \(\frac{∂y}{∂t}\). now do \(y_{tt}\), ie\(\frac{∂^2y}{∂t^2}\). just pattern match and see where you go! in terms of the x partial, you have: \(y_x =be^b(x-vt)\). do the same again to get \(y_{xx} \)